Session: May 1997 (session 2)
Frequency: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

JIVE ( at Sun Jun 8 23:39:24 METDST 1997) Minor failures
lost scan from 159 12:00:00 to 159 15:30:48

Effelsberg (dgraham at Mon Jun 9 09:24:23 METDST 1997) Minor failures
late start at 12:47 due to oil pump failure. Lost 19:00 to 19:28 and 13:47 to 14:09 due to storms

Onsala (A. Polatidis at Mon Jun 9 10:05:22 METDST 1997) Success

Jodrell (Alastairt Gunn at Mon Jun 9 10:27:30 METDST 1997) Success
No known problems.

Westerbork (tony foley at Mon Jun 9 14:50:16 METDST 1997) Success
Looks OK. Steady rain throughout the observation.

Noto (C.Trigilio at Mon Jun 9 16:38:27 METDST 1997) Success
No problem

Torun (K.Borkowski at Tue Jun 10 11:48:54 METDST 1997) Minor failures
Problems with LO. Fixed from 14:35

Calibration (Fredrik T Rantakyro at Mon Jul 21 11:10:33 METDST 1997) Minor failures
The ANTAB file for the experiment can be found on VLBEER in the MAY97 directory with the name EG15C.ANTAB. The file is ready to be used with the AIPS tasks ANTAB/APCAL. Read the ANTAB file for any specific comments regarding this experiment.

MPIfR (Scott Aaron at Mon Jun 29 11:46:09 METDST 1998) Minor failures
Correlation completed. Jodrell tape has low quality, especially on reverse passes.